# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

### v1

@@ 3.1 Graph Retriever
Existing RAG methodologies are mainly designed for plain text documents or triplets, where information retrieval occurs independently of the graph structure [13, 5, 17]. In retrieval stage, we first utilize an encoder-only language model (e.g., SentenceBERT [33]) to encode textual information, including the query $t_{q}$, text of each node $t_{n}$ and edge $t_{e}$ in graph respectively:
$\boldsymbol{q}=\text{SBERT}(t_{q})\in\mathbb{R}^{d},\boldsymbol{n}=SBERT(t_{n})\in\mathbb{R}^{d},\boldsymbol{e}=SBERT(t_{e})\in\mathbb{R}^{d}$ (1)
@@
$\mathcal{V}_{k}=\text{argtop}_{k}\big{(}\text{sim}(\boldsymbol{q},\boldsymbol{n})\big{)},\quad\mathcal{E}_{k}=\text{argtop}_{k}\big{(}\text{sim}(\boldsymbol{q},\boldsymbol{e})\big{)}$ (2)
This forms retrieve top-$k$ entities and relation sets, denoted as $\mathcal{G}_{\text{retriever}}$. Inspired by G-retriever [18], we further leverage the Prize-Collecting Steiner Tree algorithm [34] to maintain a controlled graph size, as detailed in Appendix A.
3.2 Graph Aligner
@@
3.2.1 Summarized Reasoning Chain
@@
To achieve the Aligner module’s optimization goals, our motivation is to leverage crucial reasoning chain related to query-answer pair. This is especially important for multi-hop problems, which require intermediate reasoning steps to bridge information gaps. Incorporating such reasoning significantly improves answer quality, as demonstrated by methods like Chain-of-Thoughts [35] and O1 [36]. To explicitly identify the essential nodes within graph database, we propose an LLM-based summarization technique. By designing well-crafted prompts, we harness the capabilities of strong LLMs to generate comprehensive reasoning chains that logically connect the question to the answer. For example of Table 8 of Appendix G, given a query, “What is the name of the first Harry Potter novel?”, with the LLM-generated prior, we can identify the intermediate and critical node, J.K. Rowling. We believe this can serve as an excellent label. The deeper insights provided through reasoning are subsequently utilized in the dual node and representation alignment process to train the aligner module. More details for summarization prompt are provided in Appendix G.
3.2.2 Node Alignment
@@
The Aligner module utilizes the summarized reasoning chain to identify and align relevant nodes within the graph, effectively pruning redundant nodes and edges by filtering out unrelated information at the node level. Specifically, we employ a GNN, which can be GraphTransformer [37] or GAT [38], to encode the structural information of the graph. The GNN produces node-level embeddings $\boldsymbol{n}_{g}$ based on the input graph:
$\boldsymbol{n}_{g}=\text{GNN}(\mathcal{G})\in\mathbb{R}^{|\mathcal{V}|\times d}$ (3)
@@
where, $|\mathcal{V}|$ is the number of node and $d$ is feature dim. For the reasoning text, we employ SBERT to encode the textual description into a representation $\boldsymbol{r}_{s}$ that captures its semantic meaning:
$\boldsymbol{r}_{s}=\text{SBERT}(t_{\text{reasoning}})\in\mathbb{R}^{d_{s}}$ (4)
@@
We concatenate the embedding of each node in the subgraph with the query embedding. The concatenated embeddings are then passed through an MLP module, which generates a predicted importance score for each node. Finally, these scores are transformed into probability distributions using the softmax function, producing both predict and reasoning importance scores.
$\boldsymbol{p}_{\text{predict}}=\text{Softmax}(\text{MLP}([\boldsymbol{n}_{g},\boldsymbol{q}^{\text{expanded}}]))\in\mathbb{R}^{|\mathcal{V}|},\boldsymbol{p}_{\text{reasoning}}=\text{Softmax}(\text{sim}(\boldsymbol{r}_{s},\boldsymbol{n}))\in\mathbb{R}^{|\mathcal{V}|}$ (5)
@@
where [,][,][ , ] means concat operation, $\boldsymbol{q}^{\text{expanded}}$ is the query embedding broadcasted across all nodes. To align the node distribution, we minimize the Kullback–Leibler (KL) divergence [32] between predicted probabilities $\boldsymbol{p}_{\text{predict}}$ and the probabilities $\boldsymbol{p}_{\text{reasoning}}$. The KL divergence loss for a subgraph is given by:
$\mathcal{L}_{\text{NA}}=\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\boldsymbol{p}_{\text{reasoning}}(i)\log\frac{\boldsymbol{p}_{\text{reasoning}}(i)}{\boldsymbol{p}_{\text{predict}}(i)}$ (6)
@@ 3.2.3 Representation Alignment
$\boldsymbol{r}_{g}=\text{POOL}(\boldsymbol{n}_{g})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\boldsymbol{n}_{g}(v)\in\mathbb{R}^{d}$ (7)
To unify the graph and text representations in a shared semantic space, we apply a contrastive loss with in-batch negative sampling. This loss encourages positive pairs (i.e., graph and text embeddings) to have higher similarity, while pushing apart non-matching pairs. Then, a shared-weight MLP layer is used to map $\boldsymbol{r}_{g}$ and $\boldsymbol{r}_{s}$ to dimension $d_{t}$ of LLM token embeddings:
$\boldsymbol{\hat{r}}_{s}=\text{MLP}(\boldsymbol{r}_{s})\in\mathbb{R}^{d_{t}},\boldsymbol{\hat{r}}_{g}=\text{MLP}(\boldsymbol{r}_{g})\in\mathbb{R}^{d_{t}}$ (8)
@@
$\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s})=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{j})/\tau\right)}\right]$ (9)
where, $N$ is the batch size, $(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})$ is the $i$-th positive (graph-text) pair in the batch, $\tau$ is a temperature parameter to control the sharpness. Similarly, we can obtains the loss $\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})$ from $\boldsymbol{\hat{r}}_{s}$ to $\boldsymbol{\hat{r}}_{g}$. The final representation alignment loss are obtained by:
$\mathcal{L}_{RA}=1/2(\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})+\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s}))$ (10)
To achieve an optimized Graph Aligner, we perform joint optimization of node and representation Alignment. The total loss for the Graph Aligner is defined as: $\mathcal{L}_{\text{Aligner}}=\mathcal{L}_{\text{RA}}+\mathcal{L}_{\text{NA}}$ The parameters of GNN encoder are jointly optimized using the loss $\mathcal{L}_{\text{Aligner}}$ with a specified training step, which reflects degree of alignment. We evaluate impact of the align degree in Figure 2 and the effectiveness in Section 4.4.
3.2.4 Graph Pruning and Representation Generation
We use the trained Graph Aligner to prune irrelevant nodes, and produce graph representations that are aligned with language. The Aligner performs graph pruning across the entire dataset. To start, we introduce a hyperparameter: the number of seed nodes, denoted as $n_{\text{seed}}$. For the graph’s node-level output, $\boldsymbol{n}_{\text{predict}}$, the top $n_{\text{seed}}$ nodes are selected as seed nodes. Using the first-order neighbors of these seed nodes, we expand the adjacent nodes and edges. This process yields the newly constructed subgraph, denoted as $\mathcal{G}_{\text{Aligner}}=(\mathcal{V}_{\text{Aligner}},\mathcal{E}_{\text{Aligner}})$. What is more, we also utilize the trained Aligner across the entire dataset to produce the graph representation $\boldsymbol{r}_{g}$. The aligned subgraph $\mathcal{G}_{\text{Aligner}}$ serves as the input to LLM, enabling more efficient and accurate generation results. The efficiency of pruning are evaluated in Appendix C.
3.3 Graph Generator
@@
$p_{\theta}(Y|t_{q},\mathcal{G}_{\text{Aligner}})=\prod_{i=1}^{m}p_{\theta}(y_{i}|[\text{MLP}(\boldsymbol{r}_{g}),\boldsymbol{r}_{t}],y_{<i})$ (11)
@@
To enhance efficiency of training and deployment, we fine-tune generator using parameter-efficient tuning method, like LoRA [39].
4 Experiments
We conducted experiments on the GraphQA benchmark [18], which includes ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). To ensure a fair comparison, we utilized the same retrieval results obtained via the PCST algorithm across all baselines. More details about datasets, baselines and implementation details are provided in Appendix B.
4.1 Main Results

### v2

@@ 4.1. Graph Retriever
Existing RAG methodologies are mainly designed for plain text documents or triplets, where information retrieval occurs independently of the graph structure (Fan et al., 2024; Zhu et al., 2023; Pan et al., 2024). In the retrieval stage, we first utilize an encoder-only language model (e.g., SentenceBERT (Reimers and Gurevych, 2020)) to encode textual information, including the query $t_{q}$, text of each node $t_{n}$ and edge $t_{e}$ in the graph, respectively:
(1) $\boldsymbol{q}$ $=\text{SBERT}(t_{q})\in\mathbb{R}^{d},$
(2) $\boldsymbol{n}$ $=\text{SBERT}(t_{n})\in\mathbb{R}^{d},$
(3) $\boldsymbol{e}$ $=\text{SBERT}(t_{e})\in\mathbb{R}^{d}$
@@
(4) $\mathcal{V}_{k}$ $=\operatorname*{arg\,topk}_{\begin{subarray}{c}\boldsymbol{n}\in\mathcal{V}\end{subarray}}\big(\text{sim}(\boldsymbol{q},\boldsymbol{n})\big),$
(5) $\mathcal{E}_{k}$ $=\operatorname*{arg\,topk}_{\begin{subarray}{c}\boldsymbol{e}\in\mathcal{E}\end{subarray}}\big(\text{sim}(\boldsymbol{q},\boldsymbol{e})\big)$
This forms top-$k$ entities and relation sets, denoted as $\mathcal{G}_{\text{retriever}}$. Inspired by He et al. (2024), we further leverage the Prize-Collecting Steiner Tree algorithm (PCST) (Bienstock et al., 1993) to maintain a controlled graph size.
To achieve the PCST algorithm, we assign “prize” to nodes and edges based on their similarity to a given query. Relevance is determined through the ranked sets of cosine similarity, $\mathcal{V}_{k}$ and $\mathcal{E}_{k}$, as follows:
[…]
(7) $\mathcal{G}_{\text{retriever}}=\underset{\begin{subarray}{c}S\subseteq\mathcal{G},\\ S\text{ is connected}\end{subarray}}{\arg\max}\left(\sum_{n\in\mathcal{V}_{S}}\text{prize}(n)+\sum_{e\in\mathcal{E}_{S}}\text{prize}(e)-\text{cost}(S)\right)$
4.2. Graph Aligner
@@
4.2.1. Summarized Reasoning Chain
@@
To achieve the Aligner module’s optimization goals, our motivation is to leverage a crucial reasoning chain related to the query-answer pair. This is especially important for multi-hop problems requiring intermediate reasoning steps to bridge information gaps. Incorporating such reasoning significantly improves answer quality, as demonstrated by methods like Chain-of-Thoughts (Wei et al., 2022) and O1 (Zhong et al., 2024). To explicitly identify the essential nodes within graph database, we propose an LLM-based summarization technique. By designing well-crafted prompts, we harness the capabilities of strong LLMs to generate comprehensive reasoning chains that logically connect the question to the answer. For example, Table 6 in the case study, given a query, “What is the name of the first Harry Potter novel?”, with the LLM-generated prior, we can identify the intermediate and critical node, J.K. Rowling. We believe this can serve as an excellent label. The deeper insights provided through reasoning are subsequently utilized in the dual node and representation alignment process to train the aligner module. More details for the summarization prompt are provided in experiments 5.10.
4.2.2. Node Alignment
@@
The Aligner module utilizes the summarized reasoning chain to identify and align relevant nodes within the graph, effectively pruning redundant nodes and edges by filtering out unrelated information at the node level. Specifically, we employ a GNN (e.g., GraphTransformer (Shi et al., 2020) or GAT (Veličković et al., 2018)), to encode the structural information of the graph. The GNN produces node-level embeddings $\boldsymbol{n}_{g}$ based on the input graph:
(8) $\boldsymbol{n}_{g}=\text{GNN}(\mathcal{G})\in\mathbb{R}^{|\mathcal{V}|\times d},$
@@
where $|\mathcal{V}|$ is the number of node and $d$ is feature dim. For the reasoning text, we employ SBERT to encode the textual description into a representation $\boldsymbol{r}_{s}$ that captures its semantic meaning:
(9) $\boldsymbol{r}_{s}=\text{SBERT}(t_{\text{reasoning}})\in\mathbb{R}^{d_{s}}$
@@
We concatenate the embedding of each node in the subgraph with the query embedding. The concatenated embeddings are then passed through an MLP module, which generates a predicted importance score for each node. Finally, these scores are transformed into probability distributions using the softmax function, producing importance scores for both prediction and reasoning.
(10) $\boldsymbol{p}_{\text{prediction}}$ $=\text{Softmax}(\text{MLP}([\boldsymbol{n}_{g},\boldsymbol{q}^{\text{expand}}]))\in\mathbb{R}^{|\mathcal{V}|},$
(11) $\boldsymbol{p}_{\text{reasoning}}$ $=\text{Softmax}(\text{sim}(\boldsymbol{r}_{s},\boldsymbol{n}))\in\mathbb{R}^{|\mathcal{V}|}$
@@
where $[,]$ means concat operation, $\boldsymbol{q}^{\text{expand}}$ is the query embedding broadcasted across all nodes. To align the node distribution, we minimize the Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951) between predicted probabilities $\boldsymbol{p}_{\text{prediction}}$ and the probabilities $\boldsymbol{p}_{\text{reasoning}}$. The KL divergence loss for a subgraph is given by:
(12) $\mathcal{L}_{\text{NA}}=\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\boldsymbol{p}_{\text{reasoning}}(i)\log\frac{\boldsymbol{p}_{\text{reasoning}}(i)}{\boldsymbol{p}_{\text{prediction}}(i)}$
@@ 4.2.3. Representation Alignment
(13) $\boldsymbol{r}_{g}=\text{POOL}(\boldsymbol{n}_{g})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\boldsymbol{n}_{g}(v)\in\mathbb{R}^{d}$
To unify the graph and text representations in a shared semantic space, we apply a contrastive loss with in-batch negative sampling. This loss encourages positive pairs (i.e., graph and text embeddings) to have higher similarity, while pushing apart non-matching pairs. Then, a shared-weight MLP layer is further designed to map $\boldsymbol{r}_{g}$ and $\boldsymbol{r}_{s}$ to dimension $d_{t}$ of LLM token embeddings:
(14) $\boldsymbol{\hat{r}}_{s}$ $=\text{MLP}(\boldsymbol{r}_{s})\in\mathbb{R}^{d_{t}},$
(15) $\boldsymbol{\hat{r}}_{g}$ $=\text{MLP}(\boldsymbol{r}_{g})\in\mathbb{R}^{d_{t}}$
@@
(16) $\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s})=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{j})/\tau\right)}\right]$
where, $N$ is the batch size, $(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})$ is the $i$-th positive (graph-text) pair in the batch, $\tau$ is a temperature parameter to control the sharpness. Similarly, we can obtains the loss $\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})$ from $\boldsymbol{\hat{r}}_{s}$ to $\boldsymbol{\hat{r}}_{g}$. The final representation alignment loss is obtained by:
(17) $\mathcal{L}_{RA}=\frac{1}{2}(\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})+\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s}))$
To achieve an optimized Graph Aligner, we perform joint optimization of node and representation Alignment. The total loss for the Graph Aligner is defined as:
(18) $\mathcal{L}_{\text{Aligner}}=\mathcal{L}_{\text{RA}}+\mathcal{L}_{\text{NA}}$
The parameters of GNN encoder are jointly optimized using the loss $\mathcal{L}_{\text{Aligner}}$ with a specified training step, which reflects degree of alignment. We evaluate the impact of the alignment degree in Figure 4 and the effectiveness in Section 5.7.
4.2.4. Graph Pruning and Representation Generation
We use the trained Graph Aligner to prune irrelevant nodes, and produce graph representations that are aligned with language. The Aligner performs graph pruning across the entire dataset. To start, we introduce a hyperparameter: the number of seed nodes, denoted as $n_{\text{seed}}$. For the graph’s node-level output, $\boldsymbol{n}_{\text{prediction}}$, the top $n_{\text{seed}}$ nodes are selected as seed nodes. Using the first-order neighbors of these seed nodes, we expand the adjacent nodes and edges. This process yields the newly constructed subgraph, denoted as $\mathcal{G}_{\text{Aligner}}=(\mathcal{V}_{\text{Aligner}},\mathcal{E}_{\text{Aligner}})$. What is more, we also utilize the trained Aligner across the entire dataset to produce the graph representation $\boldsymbol{r}_{g}$. The aligned subgraph $\mathcal{G}_{\text{Aligner}}$ serves as the input to LLM, enabling more efficient and accurate generation results. The efficiency of pruning is evaluated in section 5.8.
4.3. Graph Generator
@@
(19) $p_{\theta}(Y|t_{q},\mathcal{G}_{\text{Aligner}})=\prod_{i=1}^{m}p_{\theta}(y_{i}|[\text{MLP}(\boldsymbol{r}_{g}),\boldsymbol{r}_{t}],y_{<i})$
@@
To enhance efficiency of training and deployment, we fine-tune generator using parameter-efficient tuning method, like LoRA (Hu et al., 2021).
5. Experimental Settings
Table 1. Statistics of GraphQA Benchmark (He et al., 2024).
Dataset #Training #Validation #Test Task Type
ExplaGraphs 1,659 553 554 Commonsense reasoning
SceneGraphs 59,978 19,997 20,025 Visual scene understanding
WebQSP 2,826 245 1,628 Knowledge-based QA
5.1. Datasets and Metrics.
Following G-Retriever (He et al., 2024), we conducted experiments on the GraphQA benchmark (He et al., 2024), which includes ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). The dataset statistics are summarized in Table 1, with a data split of train:validation:test = 6:2:2.
  • ExplaGraphs is a dataset designed for commonsense reasoning and focuses on generating explanation graphs to predict stances in debates. It provides detailed, explicit, and commonsense-enriched graphs to evaluate arguments that either support or refute a given stance. The primary task is to determine whether an argument supports or opposes a stance.
  • SceneGraphs is a visual question answering dataset. Each graph provides detailed descriptions of objects, attributes, and relationships within an image. It is designed for tasks that involve spatial understanding and multi-step reasoning. The goal is to answer open-ended questions based on textual descriptions of scene graphs, with Accuracy used as the evaluation standard.
  • WebQSP is a large-scale, multi-hop knowledge graph question-answering dataset built on a subset of Freebase. It focuses on questions requiring multi-hop reasoning and includes facts within a two-hop neighborhood of the entities mentioned in the question. Since a single question may have multiple answers, the model’s performance is evaluated using F1, Hit@1, and Accuracy following Hu et al. (2024).
5.2. Baselines.
To assess the effectiveness of our proposed method, we compare it against four categories of baselines:
5.2.1. Inference-only.
This category includes approaches that employ frozen LLMs for question answering, using textual graph descriptions and queries as input. Different prompt strategies are considered, including:
  • Zero-shot: directly answering questions based on retrieved information,
  • Zero-CoT (Wei et al., 2022): enhances zero-shot reasoning by appending the phrase “Let’s think step by step.”, which encourages the model to explicitly generate intermediate reasoning steps.
  • CoT-BAG (Wang et al., 2024): adds “Let’s construct a graph with the nodes and edges first.” after providing the textual graph description,
  • KAPING (Baek et al., 2023): augments LLMs with knowledge graph facts at the prompt level, enabling more accurate QA without requiring model fine-tuning retrieving relevant graph triples.
  • ToG (Sun et al., [n. d.]): integrates LLMs with KGs to support deep and reliable reasoning through interactive exploration of entities and relations.
5.2.2. Raw Fine-tuning.
In this setting, following (He et al., 2024), we fine-tune the LLM via parameter-efficient methods, without applying advanced reranking techniques. Two widely used approaches are included:
  • Prompt Tuning (Lester et al., 2021): learns a set of continuous prompt embeddings that guide the frozen LLM during inference, enabling lightweight adaptation to specific tasks.
  • LoRA (Hu et al., 2021): introduces low-rank adaptation layers into transformer weights, achieving efficient fine-tuning with significantly reduced computational and memory costs.
5.2.3. Reranker-based.
These baselines employ reranking models to refine the ranking of candidate documents before passing them into the LLM. Representative methods include:
  • GTE (General Textual Embedding) (Li et al., 2023b): developed by Alibaba DAMO Academy, with two variants (gte-base, 109M parameters; gte-large, 335M parameters), trained on large-scale relevance pairs to improve retrieval accuracy,
  • BGE (BAAI General Embedding) (Xiao et al., 2024): cross-encoder models with two variants (bge-reranker-base, 278M parameters; bge-reranker-large, 560M parameters), optimized for retrieval-augmented generation, offering high accuracy at the cost of efficiency,
  • G-RAG and G-RAG-RL (Dong et al., 2024): reranker models that leverage GNNs to incorporate document connections and semantic cues from abstract meaning representation graphs for context-aware ranking.
5.2.4. GNN-based.
This category integrates GNN encoders with LLM embeddings to enhance graph reasoning, including:
  • GraphToken (Perozzi et al., 2024): encodes graph structures as explicit prompts for LLMs, improving performance on graph reasoning tasks,
  • GNP (Tian et al., 2024): a plug-and-play framework that applies a GNN encoder with cross-modality pooling to enrich LLMs with knowledge graph information,
  • G-Retriever (He et al., 2024): introduces the GraphQA benchmark and employs retrieval-augmented generation with soft prompting, improving graph-based question answering, comprehension, and reducing hallucinations,
  • GRAG (Hu et al., 2024): proposes a divide-and-conquer strategy for efficient textual subgraph retrieval and combines text and graph views within LLMs for graph context-aware generation.
5.3. Implementation Details.
We use GraphTransformer (Shi et al., 2020), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2016) as GNN encoders, and Llama-2-7b-hf, Llama-2-13b-hf and Llama-3.1-8b (Touvron et al., 2023) as generators. For the retrieval process, we use the same retrieval results obtained through the PCST algorithm across all baselines to ensure a fair comparison. Both reranker-based and GNN-based methods apply LoRA for fine-tuning, except for GraphToken, whose results are from G-Retriever (He et al., 2024). All methods are compared using the same training hyperparameters where applicable, ensuring fair comparability (e.g., GNN layers and LoRA rank). In the Aligner module, we explore two hyperparameters: alignment degree and the number of seed nodes ($n_{\text{seed}}$), with the analysis shown in Figure 7. To summarize the reasoning chain, we employ Llama-3.1-70B-Instruct (Dubey et al., 2024). The models were trained on the training set, with optimal hyperparameters and early stopping determined using the validation set. Results are reported on the test set. All experiments were run on two NVIDIA A100 GPUs (80GB each).
To implement baselines, for GNP (Tian et al., 2024), we implemented the Graph Neural Prompting module within our framework, including components such as the GNN encoder, cross-modality pooling, and domain projector. For G-RAG and G-RAG-RL (Dong et al., 2024), we adopted their ranking approach, combining cross-entropy loss with pairwise ranking loss. However, since treating documents as nodes was infeasible in our case, we instead treated entities as nodes and employed the same GNN encoder as our method. We implemented GTE (Li et al., 2023b) and BGE-reranker (Xiao et al., 2024) using their official open-source models. Nodes and triples are ranked by query similarity and, as in our method, fed into the LLM for generation. For the $\text{GRAG}_{PCST}$ model (Hu et al., 2024), we reproduced its experiments. However, to ensure a fair evaluation, we standardized the retrieval process by using the PCST method for retrieval. This allowed us to directly compare it with their graph encoder approach.
5.4. Main Results